# Jet Tagging with More-Interaction Particle Transformer

### v1

@@ Abstract
In this study, we introduce the More-Interaction Particle Transformer (MIParT), a novel deep learning neural network designed for jet tagging. This framework incorporates our own design, the More-Interaction Attention (MIA) mechanism, which increases the dimensionality of particle interaction embeddings. We tested MIParT using the top tagging and quark-gluon datasets. Our results show that MIParT not only matches the accuracy and AUC of LorentzNet but also significantly outperforms the ParT model in background rejection. Specifically, it improves background rejection by approximately 25% at a 30% signal efficiency on the top tagging dataset and by 3% on the quark-gluon dataset. Additionally, MIParT requires only 30% of the parameters and 47% of the computational complexity needed by ParT, proving that high performance can be achieved with reduced model complexity and without extensive pretraining on large datasets. The results suggest that MIParT has the potential to advance efficiency benchmarks for jet tagging and event identification in particle physics.
@@ I Introduction
Jet identification has become a key area where machine learning is applied in high-energy physics, and has made significant progress in the last few decades [1, 2]. Jets are collimated sprays of particles produced in high-energy collisions, typically from quarks, gluons, or the hadronic decay of heavy particles. The process known as jet tagging, which involves identifying the particle that initiated the jet, is complex and challenging. This complexity arises because the initial particle decays into a jet through multiple stages, increasing the number of particles within the jet and obscuring the characteristics of the initiating particle.
By analyzing the constituents of a jet, it is possible to determine the type of particle that initiated the jet. This identification is critical for revealing fundamental physical processes and discovering new particles. Initially, jet tagging relied heavily on quantum chromodynamics (QCD) theory, which provided methodologies for distinguishing between quark and gluon jets [3, 4, 5, 6, 7, 8, 9]. With the advent of machine learning, a variety of new jet tagging methods have been introduced that utilize different machine learning models to improve the breadth and accuracy of the techniques [10, 11, 12, 13, 14, 15]. Recent advances in deep learning have further refined jet tagging methods, allowing modern algorithms to effectively process large and complex datasets. These algorithms are adept at identifying subtle patterns that differentiate various types of jets, significantly improving the accuracy and efficiency of jet tagging [16, 17, 18, 19, 20]. The exceptional ability of deep learning to handle large data sets has been instrumental in these advances, leading to the discovery of new physical phenomena and deepening our understanding of particle interactions [20].
@@
In 2019, Qu et al. introduced ParticleNet [34], building on the Dynamic Graph Convolutional Neural Network (DGCNN) framework developed by Wang et al. in 2018 [35] . ParticleNet, which also treats jets as unordered sets of particles, marked significant advancements in this field. Recently, in 2022, Qu et al. further extended their contributions by developing the Particle Transformer (ParT) [36], which is based on the Transformer architecture [37]. Utilizing their own extensive dataset, JetClass, ParT underwent pre-training, significantly enhancing its capabilities. At the same time, by incorporating pairwise particle interaction inputs, it significantly improved performance on jet tagging.
@@
In this paper, we propose a new jet tagging method based on the Transformer architecture, called More-Interaction Particle Transformer (MIParT). We enhanced the algorithm of ParT by modifying the attention mechanism and increasing the embedding dimensions of the pairwise particle interaction inputs while reducing the total number of parameters and computational complexity. We tested MIParT on two widely used jet tagging benchmarks and found that it achieves improvements over existing methods.
@@ II.1 Particle Attention Block
The Particle Attention Block, a crucial element of the ParT framework, has been seamlessly integrated into our MIParT model. The architecture of this block is based on the NormFormer design, specifically using the Layer Normalization instead of the Batch Normalization. Layer Normalization optimizes normalization by adjusting each layer individually for every single sample, enhancing model stability and overall performance across diverse datasets. The architecture of the Particle Attention Block is illustrated in Fig. 3. Furthermore, in this configuration, the traditional Multi-Head Attention (MHA) is substituted by Particle-Multi-Head Attention (P-MHA). This modification allows for the incorporation of particle interaction features directly into the attention mechanism, enriching the model’s capability to capture complex particle dynamics. The P-MHA mechanism, which is key to the Particle Attention Block, is mathematically expressed as
@@ III Result and Discussion
Table 1: Summary of kinematic and particle identification variables included in the top tagging (TOP) and quark-gluon (QG) datasets. Variables present in each dataset are indicated by a star symbol ($\star$). The table lists seven kinematic variables, which describe the physical characteristics of particles relative to the jet axis, and six particle identification variables, which categorize particles based on their type and charge.
       Category Variable TOP QG
$\Delta\eta$star $\star$
$\Delta\phi$star $\star$
log $p_{\rm T}$star $\star$
       Kinematics log $E$star $\star$
$\log{p_{\rm T}}/{p_{\rm T}{\rm(jet)}}$star $\star$
$\log{E}/{E{\rm(jet)}}$star $\star$
$\Delta R$star $\star$
charge $\star$
Electron $\star$
       Particle Muon $\star$
       identification Photon $\star$
Charged Hadron $\star$
Neutral Hadron $\star$
We developed the MIParT model using the PyTorch framework [39], implemented based on the Weaver [^1] Weaver provides a streamlined yet flexible machine learning R&D framework for high energy physics, https://github.com/hqucms/weaver-core., and also referred to the implementation of particle transformer [^2] The official implementation of Particle Transformer for Jet Tagging, which includes the code and pre-trained models, https://github.com/jet-universe/particle_transformer.. The architecture of our MIParT model includes $K=5$ MI-particle attention blocks, $L=5$ particle attention blocks, and 2 class attention blocks. For particle embeddings $\mathbf{x}_{1}$, a three-layer Multi-Layer Perceptron (MLP) is used, with each layer containing 128, 512, and 64 neurons respectively. This configuration results in embeddings with a dimensionality of $D_{1}=64$. Each layer incorporates GELU as the activation function and Layer Normalization. Additionally, a three-layer, 64-channel Pointwise 1D Convolution is used for the interaction embeddings $\mathbf{U}_{1}$, performing convolutions only along the feature dimension. The $\mathbf{U}_{1}$ embeddings are further processed through a single-layer, 8-channel Pointwise 1D Convolution to generate $\mathbf{U}_{2}$, achieving a dimensionality of $D_{2}=8$. The MI-particle attention blocks implement MIA with 64 heads, while the P-MHA and Class Multi-Head Attention in the particle and class attention blocks utilize 8 heads each. A dropout rate of 0.1 is maintained in all MI-particle and particle attention blocks, with the class attention blocks being exempt from dropout. The model was trained on an NVIDIA RTX 4090 GPU, using a learning rate of 0.001 and a batch size of 256. Training was limited to 15 epochs to prevent overfitting.
@@
We evaluated the MIParT model on two widely used jet tagging benchmark datasets, top tagging [16] and quark-gluon datasets [40]. Both datasets incorporate kinematic variables as particle input features, with particle identification information included only in the quark-gluon dataset. All these input features for the two datasets are shown in Table 1.
The seven kinematic input features in the top tagging and quark-gluon datasets are:
@@
The six particle identification features in the quark-gluon dataset are:
@@
Table 2: Comparative performance of various models on the top tagging dataset. This figure illustrates the results for the MIParT model alongside those of other prominent models such as Particle Flow Network (PFN) [40], particle-level convolutional neural network (P-CNN), ParticleNet [34], Point Cloud Transformer (PCT) [42], LorentzNet [43], and ParT [36]. Metrics of other models are quoted from their published results.
Accuracy AUC $\text{Rej}_{50\%}$ $\text{Rej}_{30\%}$
PFN 0.9819 247±3 888±17
P-CNN 0.930 0.9803 201±4 759±24
PCT 0.940 0.9855 392±7 1533±101
LorentzNet 0.942 0.9868 498±18 2195±173
ParticleNet 0.940 0.9858 397±7 1615±93
ParT 0.940 0.9858 413±16 1602±81
MIParT 0.942 0.9868 505±8 2010±97
Table 3: Comparative performance of various models on the quark-gluon dataset. This table outlines the results for the MIParT model along with other significant models, including Particle Flow Network (PFN) [40], attention-based Cloud Net (ABCNet) [44], Point Cloud Transformer (PCT) [42], LorentzNet [43], and ParT [36]. Metrics of other models are cited from their published results.
Accuracy AUC $\text{Rej}_{50\%}$ $\text{Rej}_{30\%}$
PFN 0.9052 37.4±0.7
ABCNet 0.840 0.9126 42.6±0.4 118.4±1.5
PCT 0.841 0.9140 43.2±0.7 118.0±2.2
LorentzNet 0.844 0.9156 42.4±0.4 110.2±1.3
ParT 0.849 0.9203 47.9±0.5 129.5±0.9
MIParT 0.851 0.9215 49.3±0.4 133.9±1.4
TOP QG Params FLOPs
PFN 86.1k 4.62M
P-CNN 0.930 354k 15.5M
ParticleNet 0.940 370k 540M
ParT 0.940 0.849 2.14M 340M
MIParT 0.942 0.851 720.9k 160M
Figure 5: Performance metrics comparison of MIParT with other models on the top tagging dataset. This figure displays the Accuracy, AUC, ${\rm Rej}_{50\%}$, and ${\rm Rej}_{30\%}$ metrics for the MIParT model alongside Particle Flow Network (PFN) [40], particle-level convolutional neural network (P-CNN) [34], ParticleNet [34], Point Cloud Transformer (PCT) [42], LorentzNet [43], and ParT [36]. Metrics of other models are quoted from their published results. Detailed outcomes are provided in Table 2.
Figure 6: Performance metrics comparison of MIParT with other models on the quark-gluon dataset. This figure displays the Accuracy, AUC, ${\rm Rej}_{50\%}$, and ${\rm Rej}_{30\%}$ metrics for the MIParT model alongside Particle Flow Network (PFN) [40], attention-based Cloud Net (ABCNet) [44], Point Cloud Transformer (PCT) [42], LorentzNet [43], and ParT [36]. Metrics of other models are quoted from their published results. Detailed outcomes are provided in Table 2.
@@
  • AUC (Area Under the Curve): AUC provides a comprehensive measure of model performance across all classification thresholds. This metric is derived from the Receiver Operating Characteristic (ROC) curve, which plots the true positive rate (sensitivity) against the false positive rate (1-specificity) for various thresholds. This curve illustrates the trade-off between sensitivity and specificity. An AUC value can range from 0.5, which indicates no discriminatory ability (similar to random guessing), to 1.0, which represents perfect discrimination and indicates the model’s excellent ability to discriminate between classes.
@@
Top tagging is a critical task in jet tagging, which is often used in the search for new physics at the LHC. For this study, we used a top tagging dataset [16] consisting of 2 million jets, with $t\to bqq^{\prime}$ as the signal and $q/g$ as the background. This dataset only provides the energy-momentum four-vectors (kinematic features) for each particle.
@@
In Fig. 5, we showed the performance of our MIParT model compared to other popular models on the top tagging dataset. The MIParT model achieved accuracy and AUC metrics nearly identical to those of LorentzNet [43], and its ${\rm Rej}_{50\%}$ and ${\rm Rej}_{30\%}$ metrics are within the error range comparable to LorentzNet. Moreover, both MIParT and LorentzNet significantly outperformed other models, including Particle Flow Network (PFN) [40], Particle-level Convolutional Neural Network (P-CNN) [34], ParticleNet [34], Point Cloud Transformer (PCT) [42], and ParT [36], with metrics quoted from their published results. Detailed comparison results are presented in Table 2. The MIParT model significantly outperformed ParT in the top tagging benchmark, with approximately 25% better background rejection at a 30% signal efficiency. Among the models evaluated, MIParT, along with LorentzNet, ranks at the top tier, showcasing some of the most robust performances.
@@
Quark-gluon tagging is another crucial jet tagging task. Unlike the top tagging dataset, the quark-gluon dataset [40] includes not only the kinematic features of each particle, but also particle identification information. This dataset allows for a more detailed categorization of particles, including specific distinctions among electrically charged and neutral hadrons, such as pions, kaons, and protons. Additionally, like the top tagging dataset, the quark-gluon dataset contains 2 million jets, with quarks and gluons designated as the signal and background, respectively.
@@
In Fig. 6, we showed the performance of our MIParT model compared to other popular models on the quark-gluon dataset. Within this dataset, the MIParT model significantly outperforms LorentzNet across all metrics, including accuracy, AUC, ${\rm Rej}_{50\%}$, and ${\rm Rej}_{30\%}$, as well as several other models. Moreover, only the ParT model approaches the performance of our model in several metrics, but MIParT still maintains an overall lead over ParT. In comparison with other models, such as PFN [40], ABCNet [44], and PCT [42], MIParT demonstrates a substantial lead, with metrics quoted from their published results. Detailed comparison results on the quark-gluon dataset are presented in Table 3. MIParT achieves the best performance across all evaluation metrics, improving background rejection power by approximately 3% compared to ParT.
@@
In Table 4 we present the parameters, FLOPs (Floating Point Operations Per Second), and accuracy of various models on the top tagging and quark-gluon datasets. Parameters denote the number of trainable elements within a model, which indicates the capacity of the model to learn. Conversely, more parameters generally increase the complexity of the model. FLOPs measure the computational complexity required to process data through the model. Reducing the number of parameters typically reduces FLOPs, simplifying the model and making it more computationally efficient.
@@
However, reducing the number of parameters to reduce FLOPs usually results in lower accuracy. In contrast, our MIParT model has only 30% of the parameters and 47% of the FLOPs of the ParT model, significantly reducing model complexity. Despite this reduction, there is no compromise in accuracy; in fact, accuracy improves on both top tagging and quark-gluon datasets.

### v2

@@ Abstract
In this study, we introduce the More-Interaction Particle Transformer (MIParT), a novel deep learning neural network designed for jet tagging. This framework incorporates our own design, the More-Interaction Attention (MIA) mechanism, which increases the dimensionality of particle interaction embeddings. We tested MIParT using the top tagging and quark-gluon datasets. Our results show that MIParT not only matches the accuracy and AUC of LorentzNet and a series of Lorentz-equivariant methods, but also significantly outperforms the ParT model in background rejection. Specifically, it improves background rejection by approximately 25% at a 30% signal efficiency on the top tagging dataset and by 3% on the quark-gluon dataset. Additionally, MIParT requires only 30% of the parameters and 53% of the computational complexity needed by ParT, proving that high performance can be achieved with reduced model complexity. For very large datasets, we double the dimension of particle embeddings, referring to this variant as MIParT-Large (MIParT-L). We find that MIParT-L can further capitalize on the knowledge from large datasets. From a model pre-trained on the 100M JetClass dataset, the background rejection performance of the fine-tuned MIParT-L improved by 39% on the top tagging dataset and by 6% on the quark-gluon dataset, surpassing that of the fine-tuned ParT. Specifically, the background rejection of fine-tuned MIParT-L improved by an additional 2% compared to the fine-tuned ParT. The results suggest that MIParT has the potential to advance efficiency benchmarks for jet tagging and event identification in particle physics.
@@ I Introduction
Jet identification has become a key area where machine learning is applied in high-energy physics, and has made significant progress in the past few years [1, 2]. Jets are collimated sprays of particles produced in high-energy collisions, typically from quarks, gluons, or the hadronic decay of heavy particles. The process known as jet tagging, which involves identifying the particle that initiated the jet, is complex and challenging. This complexity arises because the initial particle evolves into a jet through multiple stages, increasing the number of particles within the jet and obscuring the characteristics of the initiating particle.
By analyzing the constituents of a jet, it is possible to determine the type of particle that initiated the jet. This identification is critical for revealing fundamental physical processes and discovering new particles. Initially, jet tagging relied heavily on quantum chromodynamics (QCD) theory, which provided methodologies for distinguishing between quark and gluon jets [3, 4, 5, 6, 7, 8, 9]. With the advent of machine learning, a variety of new jet tagging methods have been introduced that utilize different machine learning models to improve the breadth and accuracy of the techniques [10, 11, 12, 13, 14, 15]. Recent advances in deep learning have further refined jet tagging methods, allowing modern algorithms to effectively process large and complex datasets. These algorithms are adept at identifying subtle patterns that differentiate various types of jets, significantly improving the accuracy and efficiency of jet tagging [16, 17, 18, 19, 20]. The exceptional ability of deep learning to handle large data sets has been instrumental in these advances, leading to the discovery of new physical phenomena and deepening our understanding of particle interactions.
@@
In 2019, Qu et al. introduced ParticleNet [34], building on the Dynamic Graph Convolutional Neural Network (DGCNN) framework developed by Wang et al. in 2018 [35] . ParticleNet, which also treats jets as unordered sets of particles, marked significant advancements in this field. Recently, in 2022, Qu et al. further extended their contributions by developing the Particle Transformer (ParT) [36], which is based on the Transformer architecture [37]. By incorporating pairwise particle interaction inputs, it significantly improved performance on jet tagging. Furthermore, the introduction of a new large-scale dataset, JetClass, enables pre-training of the ParT model, which reaches even higher performance.
@@
In this paper, we propose a new jet tagging method based on the Transformer architecture, called More-Interaction Particle Transformer (MIParT). We enhanced the algorithm of ParT by modifying the attention mechanism and increasing the embedding dimensions of the pairwise particle interaction inputs while reducing the total number of parameters and computational complexity. We tested MIParT on two widely used jet tagging benchmarks and found that it achieves improvements over existing methods. Additionally, to address the challenges posed by very large datasets, we doubled the particle embedding dimensions to construct a larger model. We pre-trained this enhanced model on the 100M JetClass dataset before fine-tuning it on smaller datasets. This approach showed measurable performance gains over the fine-tuned ParT, indicating the efficacy of our modifications.
@@ II.1 Particle Attention Block
The Particle Attention Block, a crucial element of the ParT framework, has been seamlessly integrated into our MIParT model. The architecture of this block is based on the NormFormer design [39], specifically using the Layer Normalization instead of the Batch Normalization. Layer Normalization optimizes normalization by adjusting each layer individually for every single sample, enhancing model stability and overall performance across diverse datasets. The architecture of the Particle Attention Block is illustrated in Fig. 3. Furthermore, in this configuration, the traditional Multi-Head Attention (MHA) is substituted by Particle-Multi-Head Attention (P-MHA). This modification allows for the incorporation of particle interaction features directly into the attention mechanism, enriching the model’s capability to capture complex particle dynamics. The P-MHA mechanism, which is key to the Particle Attention Block, is mathematically expressed as
@@ II.3 Class Attention Block
II.4 Implementation
The architecture of our MIParT model includes $K=5$ MI-particle attention blocks, $L=5$ particle attention blocks, and 2 class attention blocks. For particle embeddings $\mathbf{x}_{1}$, a three-layer Multi-Layer Perceptron (MLP) is used, with each layer containing 128, 512, and 64 neurons respectively. This configuration results in embeddings with a dimensionality of $D_{1}=64$. Each layer incorporates GELU as the activation function and Layer Normalization. Additionally, a three-layer, 64-channel Pointwise 1D Convolution is used for the interaction embeddings $\mathbf{U}_{1}$, performing convolutions only along the feature dimension. The $\mathbf{U}_{1}$ embeddings are further processed through a single-layer, 8-channel Pointwise 1D Convolution to generate $\mathbf{U}_{2}$, achieving a dimensionality of $D_{2}=8$. The MI-particle attention blocks implement MIA with 64 heads, while the P-MHA and Class Multi-Head Attention in the particle and class attention blocks utilize 8 heads each. A dropout rate of 0.1 is maintained in all MI-particle and particle attention blocks, with the class attention blocks being exempt from dropout.
For very large datasets, increasing the embedding dimension significantly enhances model performance. Therefore, for such datasets, we double the dimension of the particle embeddings to $D_{1}=128$. This adjustment is straightforward, requiring a change in the neuron configuration of the three-layer MLP to 128, 512, and 128. Consequently, the dimensions of $\mathbf{x}$ and $\mathbf{U}$ in MIA will no longer be identical; however, this discrepancy is acceptable as long as the dimension of $\mathbf{x}$ is an integer multiple of the dimension of $\mathbf{U}$. We refer to this modified model as MIParT-Large (MIParT-L).
@@ III Result and Discussion
Table 1: Summary of kinematic and particle identification variables included in the top tagging (TOP), quark-gluon (QG) and JetClass (JC) datasets. Variables present in each dataset are indicated by a star symbol ($\star$). The table includes seven kinematic variables describing the physical characteristics of particles relative to the jet axis, six particle identification variables categorizing particles by type and charge, and four trajectory displacement features, which provide detailed information on particle trajectories.
Category Variable TOP QG JC
$\Delta\eta$star $\star$ $\star$
$\Delta\phi$star $\star$ $\star$
log $p_{\rm T}$star $\star$ $\star$
Kinematics log $E$star $\star$ $\star$
$\log{p_{\rm T}}/{p_{\rm T}{\rm(jet)}}$star $\star$ $\star$
$\log{E}/{E{\rm(jet)}}$star $\star$ $\star$
$\Delta R$star $\star$ $\star$
charge $\star$ $\star$
Electron $\star$ $\star$
Particle Muon $\star$ $\star$
identification Photon $\star$ $\star$
Charged Hadron $\star$ $\star$
Neutral Hadron $\star$ $\star$
$\tanh d_{0}$star
Trajectory $\tanh d_{z}$star
displacement $\sigma_{d_{0}}$star
$\sigma_{d_{z}}$star
We developed the MIParT model using the PyTorch framework [40], implemented based on the Weaver [^1] Weaver provides a streamlined yet flexible machine learning R&D framework for high energy physics, https://github.com/hqucms/weaver-core., and also referred to the implementation of ParT [^2] The official implementation of Particle Transformer for Jet Tagging, which includes the code and pre-trained models, https://github.com/jet-universe/particle_transformer..
@@
We initially evaluated the MIParT model on two widely used jet tagging benchmark datasets, top tagging [16] and quark-gluon datasets [41]. The model was trained on an NVIDIA RTX 4090 GPU, using a learning rate of 0.001 and a batch size of 256. Training was limited to 15 epochs to prevent overfitting. Both datasets incorporate kinematic variables as particle input features, with particle identification information included only in the quark-gluon dataset. All these input features for the two datasets are shown in Table 1.
We then pre-trained our larger model variant, MIParT-L, on the JetClass dataset containing 100M samples [36]. This model was pre-trained on dual NVIDIA RTX 3090 GPUs using a learning rate of 0.0008 and a batch size of 384, with pre-training limited to 50 epochs to avoid overfitting. After pre-training, MIParT-L was fine-tuned on the top tagging and quark-gluon datasets. It is noteworthy that the pre-training of MIParT-L on the JetClass dataset for the top tagging dataset included only kinematic features, while for the quark-gluon dataset both kinematic and particle identification features were included.
For fine-tuning, we replaced the last MLP for classification with a newly initialized MLP having two output nodes. All weights were then fine-tuned across the datasets for 20 epochs. We used a learning rate of 0.00016 for the pre-trained weights and 0.008 for the new MLP.
The seven kinematic input features are:
@@
The six particle identification features are:
@@
The four trajectory displacement features in the JetClass are:
  • $\tanh d_{0}$: hyperbolic tangent of the transverse impact parameter value;
  • $\tanh d_{z}$: hyperbolic tangent of the longitudinal impact parameter value;
  • $\sigma_{d_{0}}$: error of the measured transverse impact parameter;
  • $\sigma_{d_{z}}$: error of the measured longitudinal impact parameter.
Table 2: Comparative performance of various models on the top tagging dataset. This table displays the results for the MIParT model alongside those of other prominent models such as Particle Flow Network (PFN) [41], Particle-level Convolutional Neural Network (P-CNN), Point Cloud Transformer (PCT) [42], Clifford Group Equivariant Neural Networks (CGENN) [43], Permutation equivariant and lorentz invariant or covariant aggregator network (PELICAN) [44], Lorentz-Equivariant Geometric Algebra Transformers (L-GATr) [45], LorentzNet [46], ParticleNet [34], ParT [36]. Metrics of other models are quoted from their published results. The fine-tuned version of our model, MIParT-L f.t., is displayed at the bottom of the table for comparison with the fine-tuned ParT model, ParT f.t.
        Accuracy         AUC         $\text{Rej}_{50\%}$         $\text{Rej}_{30\%}$
              PFN         —         0.9819         247±3          888±17
              P-CNN         0.930         0.9803         201±4          759±24
              PCT         0.940         0.9855         392±7         1533±101
              CGENN         0.942         0.9869         500         2172
              PELICAN         0.9426         0.9870         —         —
              L-GATr         0.9417         0.9868         548±26         2148±106
              LorentzNet         0.942         0.9868         498±18         2195±173
              ParticleNet         0.940         0.9858         397±7         1615±93
              ParT         0.940         0.9858         413±16         1602±81
              MIParT (ours)         0.942         0.9868         505±8         2010±97
              ParT f.t.         0.944         0.9877         691±15         2766±130
              MIParT-L f.t. (ours)         0.944         0.9878         640±10         2789±133
Table 3: Comparative performance of various models on the quark-gluon dataset. This table outlines the results for the MIParT model along with other significant models, including Particle Flow Network (PFN) [41], attention-based Cloud Net (ABCNet) [47], Point Cloud Transformer (PCT) [42], LorentzNet [46], and ParT [36]. Metrics of other models are cited from their published results. The fine-tuned version of our model, MIParT-L f.t., is displayed at the bottom of the table for comparison with the fine-tuned ParT model, ParT f.t.
        Accuracy         AUC         $\text{Rej}_{50\%}$         $\text{Rej}_{30\%}$
              PFN         —         0.9052         37.4±0.7         —
              ABCNet         0.840         0.9126         42.6±0.4         118.4±1.5
              PCT         0.841         0.9140         43.2±0.7         118.0±2.2
              LorentzNet         0.844         0.9156         42.4±0.4         110.2±1.3
              ParT         0.849         0.9203         47.9±0.5         129.5±0.9
              MIParT (ours)         0.851         0.9215         49.3±0.4         133.9±1.4
              ParT f.t.         0.852         0.9230         50.6±0.2         138.7±1.3
              MIParT-L f.t. (ours)         0.853         0.9237         51.9±0.5         141.4±1.5
@@
Figure 5: Performance metrics comparison of MIParT with other models on the top tagging dataset. This figure displays the Accuracy, AUC, ${\rm Rej}_{50\%}$, and ${\rm Rej}_{30\%}$ metrics for the MIParT model alongside Particle Flow Network (PFN) [41], Particle-level Convolutional Neural Network (P-CNN), Point Cloud Transformer (PCT) [42], Clifford Group Equivariant Neural Networks (CGENN) [43], Permutation equivariant and lorentz invariant or covariant aggregator network (PELICAN) [44], Lorentz-Equivariant Geometric Algebra Transformers (L-GATr) [45], LorentzNet [46], ParticleNet [34], ParT [36]. Metrics of other models are quoted from their published results. Detailed outcomes are provided in Table 2. Bars without slashes indicate the original models without fine-tuning, while bars with slashes indicate models with fine-tuning. The gray dashed line indicates the results for MIParT, and a red dashed line shows the results for the fine-tuned MIParT-L (MIParT-L f.t.).
Figure 6: Performance metrics comparison of MIParT with other models on the quark-gluon dataset. This figure displays the Accuracy, AUC, ${\rm Rej}_{50\%}$, and ${\rm Rej}_{30\%}$ metrics for the MIParT model alongside Particle Flow Network (PFN) [41], attention-based Cloud Net (ABCNet) [47], Point Cloud Transformer (PCT) [42], LorentzNet [46], and ParT [36]. Metrics of other models are quoted from their published results. Detailed outcomes are provided in Table 2. Bars without slashes indicate the original models without fine-tuning, while bars with slashes indicate models with fine-tuning. The gray dashed line indicates the results for MIParT, and a red dashed line shows the results for the fine-tuned MIParT-L (MIParT-L f.t.).
@@
  • AUC (Area Under the Curve): AUC provides a comprehensive measure of model performance across all classification thresholds. This metric is derived from the Receiver Operating Characteristic (ROC) curve, which plots the true positive rate (sensitivity) against the false positive rate ($1-\text{specificity}$) for various thresholds. This curve illustrates the trade-off between sensitivity and specificity. An AUC value can range from 0.5, which indicates no discriminatory ability (similar to random guessing), to 1.0, which represents perfect discrimination and indicates the model’s excellent ability to discriminate between classes.
@@
Top tagging is a critical task in jet tagging, which is often used in the search for new physics at the LHC. For this study, we used a top tagging dataset [16] consisting of 2M jets, with $t\to bqq^{\prime}$ as the signal and $q/g$ as the background. This dataset only provides the energy-momentum four-vectors (kinematic features) for each particle.
@@
In Fig. 5, we showed the performance of our MIParT model compared to other popular models on the top tagging dataset. The MIParT model achieved accuracy and AUC metrics nearly identical to those of LorentzNet [46], and its ${\rm Rej}_{50\%}$ and ${\rm Rej}_{30\%}$ metrics are within the error range comparable to LorentzNet. It is noteworthy that a series of Lorentz-equivariant methods demonstrated performance similar to that of LorentzNet, such as Clifford Group Equivariant Neural Networks (CGENN) [43], Permutation equivariant and Lorentz invariant or covariant aggregator network (PELICAN) [44], Lorentz-Equivariant Geometric Algebra Transformers (L-GATr) [45]. Moreover, MIParT, LorentzNet, and several Lorentz-equivariant based models significantly outperformed other models, including Particle Flow Network (PFN) [41], Particle-level Convolutional Neural Network (P-CNN) [34], ParticleNet [34], Point Cloud Transformer (PCT) [42], and ParT [36], with metrics quoted from their published results. For the fine-tuned MIParT-L model that is pre-trained on the 100M JetClass dataset, a 39% enhancement in background rejection performance was achieved, comparable to that of the fine-tuned ParT. Detailed comparison results are presented in Table 2. The MIParT model significantly outperformed ParT in the top tagging benchmark, with approximately 25% better background rejection at a 30% signal efficiency. Among the models evaluated, MIParT, along with LorentzNet and several other Lorentz-equivariant based models, ranks at the top tier, showcasing some of the most robust performances.
         TOP          QG          Params          FLOPs
               PFN          —          86.1k          4.62M
               P-CNN          0.930          —          354k          15.5M
               ParticleNet          0.940          —          370k          540M
               ParT          0.940          0.849          2.14M          340M
               MIParT (ours)          0.942          0.851          720.9k          180M
               MIParT-L f.t. (ours)          0.944          0.853          2.38k          368M
Table 5: Comparative performance of various models on different sizes of the JetClass dataset. This table outlines the results for the MIParT-L model alongside ParticleNet [34] and ParT [36] across 2M, 10M, and 100M JetClass datasets. Metrics of other models are cited from their published results. Models trained using the full 100M training dataset are highlighted in bold text.
All classes $H\to b\bar{b}$ $H\to c\bar{c}$ $H\to gg$ $H\to 4q$ $H\to\ell\nu qq^{\prime}$ $t\to bqq^{\prime}$ $t\to b\ell\nu$ $W\to qq^{\prime}$ $Z\to qq^{\prime}$
Accuracy AUC $\text{Rej}_{50\%}$ $\text{Rej}_{50\%}$ $\text{Rej}_{50\%}$ $\text{Rej}_{50\%}$ $\text{Rej}_{99\%}$ $\text{Rej}_{50\%}$ $\text{Rej}_{99.5\%}$ $\text{Rej}_{50\%}$ $\text{Rej}_{50\%}$
ParticleNet (2 M) 0.828 0.9820 5540 1681 90 662 1654 4049 4673 260 215
ParticleNet (10 M) 0.837 0.9837 5848 2070 96 770 2350 5495 6803 307 253
ParticleNet (100 M) 0.844 0.9849 7634 2475 104 954 3339 10526 11173 347 283
ParT (2 M) 0.836 0.9834 5587 1982 93 761 1609 6061 4474 307 236
ParT (10 M) 0.850 0.9860 8734 3040 110 1274 3257 12579 8969 431 324
ParT (100 M) 0.861 0.9877 10638 4149 123 1864 5479 32787 15873 543 402
MIParT-L (2 M) 0.837 0.9836 5495 1940 95 819 1778 6192 4515 311 242
MIParT-L (10 M) 0.850 0.9861 8000 3003 112 1281 3650 16529 9852 440 336
MIParT-L (100 M) 0.861 0.9878 10753 4202 123 1927 5450 31250 16807 542 402
@@
Quark-gluon tagging is another crucial jet tagging task. Unlike the top tagging dataset, the quark-gluon dataset [41] includes not only the kinematic features of each particle, but also particle identification information. This dataset allows for a more detailed categorization of particles, including specific distinctions among electrically charged and neutral hadrons, such as pions, kaons, and protons. Additionally, like the top tagging dataset, the quark-gluon dataset contains 2M jets, with quarks and gluons designated as the signal and background, respectively.
@@
In Fig. 6, we showed the performance of our MIParT model compared to other popular models on the quark-gluon dataset. Within this dataset, the MIParT model significantly outperforms LorentzNet across all metrics, including accuracy, AUC, ${\rm Rej}_{50\%}$, and ${\rm Rej}_{30\%}$, as well as several other models. Moreover, only the ParT model approaches the performance of our model in several metrics, but MIParT still maintains an overall lead over ParT. In comparison with other models, such as PFN [41], ABCNet [47], and PCT [42], MIParT demonstrates a substantial lead, with metrics quoted from their published results. For the fine-tuned MIParT-L model pre-trained on the 100M JetClass dataset, a 6% enhancement in background rejection performance is achieved, surpassing that of the fine-tuned ParT. Detailed comparison results on the quark-gluon dataset are presented in Table 3. MIParT achieves the best performance across all evaluation metrics, improving background rejection power by approximately 3% compared to ParT. At the same time, the background rejection of the fine-tuned MIParT-L model improved by approximately 2% compared to the fine-tuned ParT.
@@
In Table 4 we present the parameters, FLOPs (Floating Point Operations Per Second), and accuracy of various models on the top tagging and quark-gluon datasets. Parameters denote the number of trainable elements within a model, which indicates its capacity to learn. Conversely, more parameters generally increase the complexity of the model. FLOPs measure the computational complexity required to process data through the model. Reducing the number of parameters typically reduces FLOPs, simplifying the model and making it more computationally efficient.
@@
However, reducing the number of parameters to reduce FLOPs usually results in lower accuracy. In contrast, our MIParT model has only 30% of the parameters and 53% of the FLOPs of the ParT model, significantly reducing model complexity. Despite this reduction, there is no compromise in accuracy; in fact, accuracy improves on both top tagging and quark-gluon datasets. For the fine-tuned version of MIParT-L, the parameters and FLOPs are comparable to those of the ParT model, but with a slight improvement in accuracy.
In Table 5, we present comparative performance of various models on different sizes of the JetClass dataset. we displaced the results for the MIParT-L model alongside ParticleNet [34] and ParT [36] across 2M, 10M, and 100M JetClass datasets. We observe that as the dataset size increases, the performance of the models improves. Specifically, MIParT-L and ParT exhibit nearly identical effectiveness on very large datasets, surpassing that of ParticleNet.